Andersonly:

( w -1.5 )(w -1.5 ) = 352.25 find the square root of each side and solve for w. Show your work and both solutions. subtract 350 from both sides, and then factor the equation for the first one one then factor if we change the equation so that it has a perfect square trinomial for the second one.

Andersonly:

\[( w -1.5 )(w -1.5 ) = 352.25 \]

snowflake0531:

what is this Show your work and both solutions. subtract 350 from both sides, and then factor the equation for the first one one then factor if we change the equation so that it has a perfect square trinomial for the second one.

Andersonly:

yeah but how should I factor it

Andersonly:

sorry its my first time on something like this

snowflake0531:

so, did you know that (w−1.5)(w−1.5) can be written as (w-1.5)^2

Andersonly:

no I didn't

Andersonly:

thank you tho

snowflake0531:

okay, now you do, right for example 3 times 3 times 3, can be written as 3^3 4x times itself 3 times, would be written as 4x^3 so, here, you can just write (w-1.5)^2 =352.25 do you know how to get that square thing out?

Andersonly:

hmm not really :v

snowflake0531:

okay, well if one side is squared, you can just square root it to get it off sqrt((w-21.5)^2) = sqrt 352.25 so, w-1.5 = sqrt 352.25 can you put into your calculator, sqrt352.25

Andersonly:

ok Ill do that

Andersonly:

324

Andersonly:

yes?